# Efficient Prior Calibration From Indirect Data

### v1

@@ Abstract
Keywords: inverse problems, generative models, prior learning, operator learning, differential equations
@@ 1.1 Set-Up
At the heart of all these loss functions is a matching of distributions. In practice both $\nu$ and the pushforward of $\mu_{0}$ will only be available empirically and so it is necessary that the divergence $\mathsf{d}_{1}$ can be readily evaluated on empirical measures. Empiricalized versions of the functionals $\mathsf{J}_{2},\mathsf{J}_{3}$ will form the basis of computational methodology proposed in this paper. The mapping $\phi^{*}(\alpha)$ will also be learned using minimization of a loss function, involving matching of distributions and evaluated empirically. The functional $\mathsf{J}_{1}$ provides a theoretical underpinning of our approach; and in the case $N=1$ will be linked, in the empirical setting, to Bayes Theorem.
@@ 1.2 Contributions and Outline
In Section 2 we describe the efficient residual-based approach to operator learning that we adopt in this paper, addressing contribution 2. Section 3 introduces specific divergences for definition of Functionals 1–3, and for the residual-based learning, addressing contributions 1. and 3. ; we also address contribution 4, in Theorem 3.2. Section 4 discusses algorithmic details, giving further detail on contributions 1. and 3. whilst Section 5 implements the algorithms on the Darcy flow problem, contribution 5. We conclude and discuss future works in Section 6.
@@ 1.3.1 Learning Priors
The task of selecting a best prior has received much attention in the Bayesian statistics community, with many objectives and ideals in mind [32]. Certain efforts concentrate on the careful formulation of uninformative priors [33, 5, 6]. Others focus on mathematical tractability through conjugacy [51] or eliciting priors from domain experts [48]. Some see the prior as an opportunity to share information from observations with possibly different underlying parameters, that are assumed to be drawn from the same distribution. Taking this point of view, are methods related to hierarchical Bayes [28], empirical Bayes [54, 55] and parametric empirical Bayes [15, 44]. In many ways, the idea of using a set of data to explicitly target an unconditional distribution is the basis of many modern generative modeling methods in the field of machine learning (ML). These include score-based models [64], diffusion models [62, 30], variational auto-encoders [36], energy-based models  [63, 67], normalizing flows [49], gradient-flows [10], and more. Data-based unconditional distributions like these have been shown to be not only expressive generative models, but also powerful priors for Bayesian inversion [25, 11]. These ideas appear in a number of recent papers focussed on inverse-problems [50, 3, 2]. A key difference in our work and the basis for our contribution is that, while above works focus on learning (and using) priors from which unconditional samples are available, our methodology here extends this idea to learn priors from indirect data. This idea of indirect knowledge of priors is exploited in [27] for linear operators where the set of Bayesian inverse problems defined individually by (1), for each $n$, is solved in the case where ${\mathcal{G}}(\cdot)=A\cdot$ for some linear operator $A$; the collection of inverse problems is used to both learn prior information and learn the dependence of each individual posterior on data.
@@ 1.3.2 Surrogate Modeling and Operator Learning
In many engineering applications related to design optimization, parameter inversion, and forward uncertainty quantification tasks, it is necessary to evaluate numerical models many times. When using a classical numerical scheme, there is no information carry-over from one numerical solve to the next. Hence, there is room for improvement in the form of somehow interpolating information from the one numerical solution to the next. The idea of replacing computer code with a cheap-to-evaluate statistical interpolation model is well explored [56, 34]. Such methods have found their place as viable model order reduction techniques for multi-query problems. Modern advances have now begun to pose the task of surrogate modeling directly in function-space, resulting in the field of operator learning [8, 42, 38, 39]. These methods are based on gathering input-output datasets of PDE parameters and PDE solutions obtain via classical numerical schemes, such as finite element models (FEM) and spectral methods. Physics informed surrogate modeling attempts to directly incorporate PDE information into the learning task [53, 73, 40]. The methods may then be data-less or semi-data-informed. Certain classes of methods also pose prior distributions over PDE parameters and attempt to learn a surrogate trained on random draws from that prior [68, 69].
@@ 1.3.3 Pushforwards and Wasserstein Losses
Minimization of regularized loss functions over the space of probability measures, as exemplified by Functional 1, is central to modern computational statistics and machine learning: it lies at the heart of variational inference [71] and in many other emerging inference problems for probability measures [18, 41]. In statistical inference and ML, there are many learning objective functions to make use of. A well known method is maximizing the marginal likelihood. However, the task we are interested in for this work, is distributional learning. Hence, we must use a statistical divergence between measures. A common choice is the KL divergence. However this is not useful when both measures being compared are empirical, due to the non-overlapping support of sampled Diracs, as is the case in this paper.
Noting this restriction on suitable divergences we proceed to identify appropriate choices. The (MMD) maximum mean divergence [61, 23] and (closely related) energy distances [65, 59] are one possible class metrics that could be used. In this work, we focus on the use of computationally tractable optimal transport based metrics on the space of measures; in particular we use the sliced-Wasserstein metric [9]. Many other works have explored the use of these optimal transport (OT) based metrics for ML inference tasks [37, 21, 45, 47, 41]. There is also work on combining pushforward measures and Bayesian inference in [14, 13, 43, 24]. Solving inverse problems with Wasserstein loss is also a partially explored topic in [1] and conditional flow matching [17]. Some works have also looked at Wasserstein metrics between pushforward measures [57]. Wasserstein and sliced-Wasserstein metrics have found use in approximate Bayesian computation (ABC) [66, 4, 46] where the recovered measure have been shown to converge to the Bayesian posterior. Parameter estimation with Wasserstein metrics in a purer form is explored in [7]. In discussing OT based learning objective, it is important to mention entropy regularization, as in the Sinkhorn algorithm [19]. In these works, entropy regularization is put on the finite-dimensional observational space. We will see that entropy regularization comes up differently in our proposed methodology.
@@ 1.4 Notation
(3a) $\nabla\cdot(z\nabla u)+f$ $=0,\quad\forall x\in\text{D},$
@@
Here $z$ denotes permeability and $u$ the piezometric head. The mapping $z\mapsto u$ may be viewed as mapping $F^{\dagger}:Z\to H^{1}_{0}(D),$ for appropriately defined $Z$. To be concrete we fill focus on this setting, and the problem of determining a prior on $z$ from noisy linear functionals of $u$ defined by mapping $g:H^{1}_{0}(D)\to{\mathbb{R}}^{d_{y}}\cong\bigl{(}H^{-1}(D)\bigr{)}^{d_{y}}.$ The reader will readily see that the ideas in the paper apply more generally, and that consideration of the Darcy problem is simply for expository purposes.
@@ 2 Residual-Based Neural Operator
This section is devoted to defining $\phi^{*}(\alpha)$ which appears in Functional 3. Recall that $\phi$ are the parameters of the neural network surrogate PDE model, and that their optimization depends on the underlying input measure on which the surrogate needs to be accurate; for this reason they depend on $\alpha$ the set of parameters characterizing the prior we seek to determine. To define function $\phi^{*}(\alpha)$ we proceed as follows. Let $F^{\phi}:Z\to H^{1}_{0}(D)$ be a parametric family of maps approximating $F^{\dagger}$. Now define residual operator $R:Z\times H_{0}^{1}(D)\to H^{-1}(D)$. Intuitively, we can express our PDE as $R(z,u)=0.$ In the case of the Darcy equation (3) we may write
@@
here $\mathsf{d}_{2}$ is a divergence term between probability measures on $\mathbb{R}^{d_{o}}$ and $\delta_{0}$ is the Dirac measure at zero. Minimizing $\mathsf{J}_{4}(\cdot;\alpha)$ for given $\alpha$ determines a residual-based approximation of $F^{\dagger}$, accurate with respect to $(T^{\alpha})_{\#}\mu_{0}$. Using this expression shows that Functional 3, Functional 4 define a bilevel optimization scheme [60]. This scheme is the heart of our proposed methodology. Using the bilevel approach avoids balancing the contributions of Functional 3 and Functional 4 that arise from an additive approach. A similar bilevel optimization scheme is employed in [72], for similar reasons, to solve a different problem. In the next section we provide specific examples $\mathsf{d}_{1}$ and $\mathsf{d}_{2}$ and discuss some of their properties.
@@ 3 Choice of Divergences
In order to instantiate the bilevel optimzation scheme defined by Functional 3, Functional 4 to obtain an implementable algorithm, we now define specific choices of the divergence terms $\mathsf{d}_{1}$ and $\mathsf{d}_{2}$ (Subsection 3.1) and discuss empirical approximation of the input measures (Subsection 3.2) required to evaluate these divergences in practice. And, to further establish a context for our work on learning priors, we make a connection to Bayesian inversion in the setting of equation (1) when $N=1$ (Subsection 3.3.)
@@ 3.3 Connection to Bayes Theorem
(13a) ${\mathrm{d}}\mu^{y}(A)$ $=\frac{1}{\mathcal{Z}}\int_{A}\exp\left(-\frac{1}{2}\|y-{\mathcal{G}}(z)\|^{2}_{\Gamma}\right){{\mathrm{d}}\mu^{\mathrm{prior}}(z)},$
@@
(15) $\mathsf{W}_{2,\Gamma}^{2}(\delta_{y},{\mathcal{G}}_{\#}\mu)=\mathbb{E}_{z\sim\mu}\|y-{\mathcal{G}}(z)\|_{\Gamma}^{2}=\int_{\mathbb{R}^{d}}\|y-{\mathcal{G}}(z)\|_{\Gamma}^{2}{\mathrm{d}}\mu(z).$
@@
Proof.
See Appendix A.4. ∎
Proof of Theorem 3.2.
By equation (15) and Lemmas 3.5 and 3.6 it follows that the loss function (12) can be written as
@@
where $C_{1}$ is independent of $\mu.$ With the choice $\mathcal{H}(\mu):=D_{\mathrm{KL}}(\mu||\mu^{\mathrm{prior}})$ we see that $\mathsf{J}(\mu)=D_{\mathrm{KL}}(\mu||\mu^{y})+C_{2}$ where $C_{2}$ is independent of $\mu$, and is determined by $C_{1}$ and $\mathcal{Z}.$ The result follows since $D_{\mathrm{KL}}(\cdot||\mu^{y})$ is minimized at $\mu^{y}.$
@@
In order to use Lemmas 3.5 and 3.6 in the proof of Theorem 3.2, we assume the pushforward measure $\eta*({\mathcal{G}}_{\#}\mu)$ has $\|\cdot\|_{\Gamma}^{2}$-moment finite as stated in Theorem 3.2. This will follow from second moments of ${\mathcal{G}}_{\#}\mu.$ Such results can often be established using the Fernique theorem; see for example [16, 20].
@@ 4 Algorithms
4:        Sample $z^{(i)}\sim\mu^{\alpha_{t-1}}$
5:        Sample $\varepsilon_{\eta}^{(i)}\sim\eta$
6:        Compute ${y^{\prime}}^{(i)}=g\circ F^{\dagger}(z^{(i)})+\varepsilon_{\eta}^{(i)}$
7:        Sample $y^{(i)}\sim\nu^{N}$
8:     Compute $\mathsf{J}_{2}(\alpha_{t-1})$ from the $N_{s}$ samples $\{y^{(i)},{y^{\prime}}^{(i)}\}_{i=1}^{N_{s}}$.
9:     $\alpha_{t}=\text{OPTIMISER}(\alpha_{t-1},\mathsf{J}_{2}(\alpha_{t-1}))$
10:  return  $\alpha^{\star}\leftarrow\alpha_{T}$.
@@
1:  Initialize $\alpha_{0},\phi_{0}$, $T$ (number of iterations of outer-loop $\alpha$ updates), $L$ (number of inner-loop $\phi$ updates), $N_{s}$ (number of samples for $\mathsf{J}_{3}$), $N_{r}$ (number of samples for $\mathsf{J}_{4}$)
@@
5:           Sample $z^{(j)}\sim\mu^{\alpha_{t-1}}$
6:           Compute $r^{(j)}=\mathcal{O}^{\phi_{l-1}(\alpha_{t-1})}(z^{(j)})$
7:        Compute $\mathsf{J}_{4}(\alpha_{t-1},\phi_{l-1}(\alpha_{t-1}))$ from the $N_{r}$ samples $\{r^{(j)}\}_{j=1}^{N_{r}}$.
8:        $\phi_{l}(\alpha_{t-1})=\text{OPTIMISER}(\phi_{l-1}(\alpha_{t-1}),\mathsf{J}_{4}(\alpha_{t-1},\phi_{l-1}(\alpha_{t-1}))$
9:     $\phi^{\star}(\alpha_{t-1})\leftarrow\phi_{L}(\alpha_{t-1})$
10:     for $i=1,\ldots,N_{s}$ do
11:        Sample $z^{(i)}\sim\mu^{\alpha_{t-1}}$
12:        Sample $\varepsilon_{\eta}^{(i)}\sim\eta$
13:        Compute ${y^{\prime}}^{(i)}=g\circ F^{\phi^{\star}(\alpha_{t-1})}(z^{(i)})+\varepsilon_{\eta}^{(i)}$
14:        Sample $y^{(i)}\sim\nu^{N}$
15:     Compute $\mathsf{J}_{3}(\alpha_{t-1},\phi^{\star}(\alpha_{t-1}))$ from the $N_{s}$ samples $\{y^{(i)},{y^{\prime}}^{(i)}\}_{i=1}^{N_{s}}$.
16:     $\alpha_{t}=\text{OPTIMISER}(\alpha_{t-1},\mathsf{J}_{3}(\alpha_{t-1},\phi^{\star}(\alpha_{t-1})))$
17:  return  $\{\alpha^{\star}\leftarrow\alpha_{T},\phi^{\star}(\alpha^{\star})\leftarrow\phi^{\star}(\alpha_{T})\}$.

### v2

@@ Abstract
keywords:
inverse problems, generative models, prior learning, operator learning, differential equations
{MSCcodes}
68T37, 65N30, 62F15, 35R30
@@ 1.1 Set-Up
At the heart of all these loss functions is a matching of distributions. In practice both $\nu$ and the pushforward of $\mu_{0}$ will only be available empirically and so it is necessary that the divergence $\mathsf{d}_{1}$ can be readily evaluated on empirical measures. Empiricalized versions of the functionals $\mathsf{J}_{2},\mathsf{J}_{3}$ will form the basis of computational methodology proposed in this paper. The mapping $\phi^{*}(\alpha)$ will also be learned using minimization of a loss function, involving matching of distributions and evaluated empirically. The functional $\mathsf{J}_{1}$ provides a theoretical underpinning of our approach; and in the case $N=1$ will be linked, in the empirical setting, to Bayes Theorem. The efficiency of the proposed methodology is due to: (i) the replacement of costly PDE simulations with evaluations of a concurrently learned surrogate model trained through readily computable PDE residuals; (ii) as will be mentioned in Remark 3.4, the proposed method scales more favorably than alternative Bayesian approaches due to the challenges of high dimensional posterior sampling.
@@ 1.2 Contributions and Outline
In Section 2 we describe the efficient residual-based approach to operator learning that we adopt in this paper, addressing contribution 2. Section 3 introduces specific divergences for definition of Functionals 1–3, and for the residual-based learning, addressing contributions 1 and 3; we also address contribution 4 in Theorem 3.5. Section 4 discusses algorithmic details, giving further detail on contributions 1 and 3 whilst Section 5 implements the algorithms on the Darcy flow problem, contribution 5. We conclude and discuss future works in Section 6.
@@ 1.3.1 Learning Priors
The task of selecting a best prior has received much attention in the Bayesian statistics community, with many objectives and ideals in mind [34]. Certain efforts concentrate on the careful formulation of uninformative priors [35, 5, 6]. Others focus on mathematical tractability through conjugacy [56] or eliciting priors from domain experts [53]. Some see the prior as an opportunity to share information from observations with possibly different underlying parameters, that are assumed to be drawn from the same distribution. Taking this point of view, are methods related to hierarchical Bayes [29], empirical Bayes [59, 60] and parametric empirical Bayes [15, 49]. In many ways, the idea of using a set of data to explicitly target an unconditional distribution is the basis of many modern generative modeling methods in the field of machine learning (ML). These include score-based models [69], diffusion models [67, 31], variational auto-encoders [38], energy-based models  [68, 72], normalizing flows [54], gradient-flows [10], and more. Data-based unconditional distributions like these have been shown to be not only expressive generative models, but also powerful priors for Bayesian inversion [26, 11]. These ideas appear in a number of recent papers focussed on inverse-problems [55, 3, 2]. A key difference in our work and the basis for our contribution is that, while above works focus on learning (and using) priors from which unconditional samples are available, our methodology here extends this idea to learn priors from indirect data. This idea of indirect knowledge of priors and data is exploited in [20] and also in [28] for linear operators where the set of Bayesian inverse problems defined individually by (1), for each $n$, is solved in the case where ${\mathcal{G}}(\cdot)=A\cdot$ for some linear operator $A$; the collection of inverse problems is used to both learn prior information and learn the dependence of each individual posterior on data.
@@ 1.3.2 Surrogate Modeling and Operator Learning
In many engineering applications related to design optimization, parameter inversion, and forward uncertainty quantification tasks, it is necessary to evaluate numerical models many times. When using a classical numerical scheme, there is no information carry-over from one numerical solve operation to the next. Hence, there is room for improvement in the form of somehow interpolating information from the one numerical solution to the next. The idea of replacing computer code with a cheap-to-evaluate statistical interpolation model is well explored [61, 36]. Such methods have found their place as viable model order reduction techniques for multi-query problems. Modern advances have now begun to pose the task of surrogate modeling directly in function-space, resulting in the field of operator learning [8, 47, 40, 44]. These methods are based on gathering input-output datasets of PDE parameters and PDE solutions obtain via classical numerical schemes, such as finite element models (FEM) and spectral methods. Physics informed surrogate modeling attempts to directly incorporate PDE information into the learning task [58, 78, 45]. The methods may then be data-less or semi-data-informed. Like PINO [45] our optimization objective balances a data-driven loss with a PDE-enforcing regularizer. Unlike PINO, however, our data-driven loss is not related to operator learning; operator learning is introduced purely through the PDE-enforcing regularizer. Also, unlike PINO, which adds the loss and the regularizer, we adopt a bilevel optimization strategy because of issues related to balancing the two terms. Finally we use the variational form of the PDE to define the regularizer, demonstrating that the FNO operator interacts well with FEM-interpolation to enable derivative-based optimization. Certain classes of methods also pose prior distributions over PDE parameters and attempt to learn a surrogate trained on random draws from that prior [73, 74].
@@ 1.3.3 Pushforwards and Wasserstein Losses
Minimization of regularized loss functions over the space of probability measures, as exemplified by Functional 1, is central to modern computational statistics and machine learning: it lies at the heart of variational inference [76] and in many other emerging inference problems for probability measures [18, 46, 43, 41, 42]. In statistical inference and ML, there are many learning objective functions to make use of. A well known method is maximizing the marginal likelihood. However, the task we are interested in for this work is distributional learning. Hence, we must use a statistical divergence between measures. A common choice is the KL divergence. However this is not useful when both measures being compared are empirical, due to the non-overlapping support of sampled Diracs, as is the case in this paper (the $D_{\mathrm{KL}}$ is either 00 or $\infty$).
Noting this restriction on suitable divergences we proceed to identify appropriate choices. The (MMD) maximum mean divergence [66, 24] and (closely related) energy distances [70, 64] are one possible class of metrics that could be used. In this work, we focus on the use of computationally tractable optimal transport based metrics on the space of measures; in particular we use the sliced-Wasserstein metric [9]. Many other works have explored the use of these optimal transport (OT) based metrics for ML inference tasks [39, 22, 50, 52, 46]. There is also work on combining pushforward measures and Bayesian inference in [14, 13, 48, 25]. Solving inverse problems with Wasserstein loss is also a partially explored topic in [1] and conditional flow matching [17]. Some works have also looked at Wasserstein metrics between pushforward measures [62]. Wasserstein and sliced-Wasserstein metrics have found use in approximate Bayesian computation (ABC) [71, 4, 51] where the recovered measure have been shown to converge to the Bayesian posterior. Parameter estimation with Wasserstein metrics in a purer form is explored in [7]. In discussing OT based learning objective, it is important to mention entropy regularization, as in the Sinkhorn algorithm [19]. In these works, entropy regularization is put on the finite-dimensional observational space. We will see that entropy regularization comes up differently in our proposed methodology.
@@ 1.4 Notation
(3a) $\nabla\cdot(z\nabla u)+f$ $=0,\quad\forall x\in D,$
@@
Here $z$ denotes permeability and $u$ the piezometric head. The mapping $z\mapsto u$ may be viewed as mapping $F^{\dagger}:Z\to H^{1}_{0}(D),$ for appropriately defined $Z$. To be concrete we will focus on this setting, and the problem of determining a prior on $z$ from noisy linear functionals of $u$ defined by mapping $g:H^{1}_{0}(D)\to{\mathbb{R}}^{d_{y}}\cong\bigl{(}H^{-1}(D)\bigr{)}^{d_{y}}.$ The reader will readily see that the ideas in the paper apply more generally, and that consideration of the Darcy problem is simply for expository purposes.
@@ 2 Residual-Based Neural Operator
This section is devoted to defining $\phi^{*}(\alpha)$ which appears in Functional 3. Recall that $\phi$ are the parameters of the neural network surrogate PDE model, and that their optimization depends on the underlying input measure on which the surrogate needs to be accurate; for this reason they depend on $\alpha$, i.e. the set of parameters characterizing the prior we want to learn. To define function $\phi^{*}(\alpha)$ we proceed as follows. Let $F^{\phi}:Z\to H^{1}_{0}(D)$ be a parametric family of maps approximating $F^{\dagger}$. Now define residual operator $R:Z\times H_{0}^{1}(D)\to H^{-1}(D)$. Intuitively, we can express our PDE as $R(z,u)=0.$ In the case of the Darcy equation (3) we may write
@@
here $\mathsf{d}_{2}$ is a divergence term between probability measures on $\mathbb{R}^{d_{o}}$ and $\delta_{0}$ is the Dirac measure at zero. Minimizing $\mathsf{J}_{4}(\cdot;\alpha)$ for given $\alpha$ determines a residual-based approximation of $F^{\dagger}$, accurate with respect to $(T^{\alpha})_{\#}\mu_{0}$. Using this expression shows that Functional 3 and Functional 4 define a bilevel optimization scheme [65, 32]. This scheme is the heart of our proposed methodology. Using the bilevel approach avoids balancing the contributions of Functional 3 and Functional 4 that arise from an additive approach. A similar bilevel optimization scheme is employed in [77], for similar reasons, to solve a different problem. In the next section we provide specific examples $\mathsf{d}_{1}$ and $\mathsf{d}_{2}$ and discuss some of their properties.
@@ 3 Choice of Divergences
In order to instantiate the bilevel optimization scheme defined by Functional 3, Functional 4 to obtain an implementable algorithm, we now define specific choices of the divergence terms $\mathsf{d}_{1}$ and $\mathsf{d}_{2}$ (Subsection 3.1) and discuss empirical approximation of the input measures (Subsection 3.2) required to evaluate these divergences in practice. And, to further establish a context for our work on learning priors, we make a connection to Bayesian inversion in the setting of equation (1) when $N=1$ (Subsection 3.3.)
@@ 3.1 Divergences
Remark 3.3.
We highlight that using the squared Wasserstein-2 metric between a Dirac at zero and an abitrary measure reduces to computing an expected squared 2-norm of the samples drawn from that measure (Lemma 3.7). Hence, (10b) reduces to computing
$\mathsf{J}_{4}(\phi;\alpha)=\mathbb{E}_{z\sim(T^{\alpha})_{\,\#}\mu_{0}}\|\mathcal{O}^{\phi}(z)\|^{2}_{2}.$
Framing (10b) this way gives a different interpretation to the commonly used physics-informed ML loss function.
@@ 3.2 Empiricalization
Remark 3.4.
In contrast to empirical/hierarchical Bayesian approaches for similar problems, the proposed methodology side-steps the sampling of a challenging posterior distribution with dimension that grows like $O(NM)$, where $N$ is the number of physical systems from which we have data, and $M$ is the dimensionality of the parameters, $z^{(n)}$, of the individual physical systems. Instead, this is replaced with the straightforward empiricalization of a pushforward measure with $N_{s}$ samples of dimension $M$.
In the numerical examples explored in Section 5, for 1D Darcy $M=20$, for 2D Darcy $M=400$, these will be the number of bases parametrizing the permiability field in the two setups respectively.
@@ 3.3 Connection to Bayes Theorem
(13a) $\mu^{y}(A)$ $=\frac{1}{\mathcal{Z}}\int_{A}\exp\left(-\frac{1}{2}\|y-{\mathcal{G}}(z)\|^{2}_{\Gamma}\right){{\mathrm{d}}\mu^{\mathrm{prior}}(z)},$
@@
(15) $\mathsf{W}_{2,\Gamma}^{2}(\delta_{y},({\mathcal{G}})_{\#}\mu)=\mathbb{E}_{z\sim\mu}\|y-{\mathcal{G}}(z)\|_{\Gamma}^{2}=\int_{\mathbb{R}^{d}}\|y-{\mathcal{G}}(z)\|_{\Gamma}^{2}{\mathrm{d}}\mu(z).$
@@
Proof 3.12.
See Appendix A.4.
Proof 3.13 (Proof of Theorem 3.5).
By equation (15) and Lemmas 3.9 and 3.11 it follows that the loss function (12) can be written as
@@
where $C_{1}$ is independent of $\mu.$ With the choice $\mathcal{H}(\mu):=D_{\mathrm{KL}}(\mu||\mu^{\mathrm{prior}})$ we see that $\mathsf{J}(\mu)=D_{\mathrm{KL}}(\mu||\mu^{y})+C_{2}$, where $C_{2}$ is independent of $\mu$, and is determined by $C_{1}$ and $\mathcal{Z}.$ The result follows since $D_{\mathrm{KL}}(\cdot||\mu^{y})$ is minimized at $\mu^{y}.$
@@
In order to use Lemmas 3.9 and 3.11 in the proof of Theorem 3.5, we assume the pushforward measure $\eta*({\mathcal{G}}_{\#}\mu)$ has $\|\cdot\|_{\Gamma}^{2}$-moment finite as stated in Theorem 3.5. This will follow from second moments of ${\mathcal{G}}_{\#}\mu.$ Such results can often be established using the Fernique theorem; see for example [16, 21].
@@ 4 Algorithms
4:        Sample $\varepsilon_{0}^{(i)}\sim\mu_{0}$
5:        Sample $z^{(i)}\sim T^{\alpha_{t-1}}(\varepsilon_{0}^{(i)})$
6:        Sample $\varepsilon_{\eta}^{(i)}\sim\eta$
7:        Compute ${y^{\prime}}^{(i)}=g\circ F^{\dagger}(z^{(i)})+\varepsilon_{\eta}^{(i)}$
8:        Sample $y^{(i)}\sim\nu^{N}$
9:     Compute $\mathsf{J}_{2}(\alpha_{t-1})$ from the $N_{s}$ samples $\{y^{(i)},{y^{\prime}}^{(i)}\}_{i=1}^{N_{s}}$.
10:     $\alpha_{t}=\text{OPTIMISER}(\alpha_{t-1},\mathsf{J}_{2}(\alpha_{t-1}))$
11:  return  $\alpha^{\star}\leftarrow\alpha_{T}$.
@@
1:  Initialize $\alpha_{0},\phi_{0}$, $T$ (number of outer-loop $\alpha$ updates), $L$ (number of inner-loop $\phi$ updates), $N_{s}$ (number of samples for $\mathsf{J}_{3}$), $N_{r}$ (number of samples for $\mathsf{J}_{4}$)
@@
5:           Sample $\varepsilon_{0}^{(j)}\sim\mu_{0}$
6:           Sample $z^{(j)}\sim T^{\alpha_{t-1}}(\varepsilon_{0}^{(i)})$
7:           Compute $r^{(j)}=\mathcal{O}^{\phi_{l-1}(\alpha_{t-1})}(z^{(j)})$
8:        Compute $\mathsf{J}_{4}(\alpha_{t-1},\phi_{l-1}(\alpha_{t-1}))$ from the $N_{r}$ samples $\{r^{(j)}\}_{j=1}^{N_{r}}$.
9:        $\phi_{l}(\alpha_{t-1})=\text{OPTIMISER}(\phi_{l-1}(\alpha_{t-1}),\mathsf{J}_{4}(\alpha_{t-1},\phi_{l-1}(\alpha_{t-1}))$
10:     $\phi^{\star}(\alpha_{t-1})\leftarrow\phi_{L}(\alpha_{t-1})$
11:     for $i=1,\ldots,N_{s}$ do
12:        Sample $\varepsilon_{0}^{(i)}\sim\mu_{0}$
13:        Sample $z^{(i)}\sim T^{\alpha_{t-1}}(\varepsilon_{0}^{(i)})$
14:        Sample $\varepsilon_{\eta}^{(i)}\sim\eta$
15:        Compute ${y^{\prime}}^{(i)}=g\circ F^{\phi^{\star}(\alpha_{t-1})}(z^{(i)})+\varepsilon_{\eta}^{(i)}$
16:        Sample $y^{(i)}\sim\nu^{N}$
17:     Compute $\mathsf{J}_{3}(\alpha_{t-1},\phi^{\star}(\alpha_{t-1}))$ from the $N_{s}$ samples $\{y^{(i)},{y^{\prime}}^{(i)}\}_{i=1}^{N_{s}}$.
18:     $\alpha_{t}=\text{OPTIMISER}(\alpha_{t-1},\mathsf{J}_{3}(\alpha_{t-1},\phi^{\star}(\alpha_{t-1})))$
19:  return  $\{\alpha^{\star}\leftarrow\alpha_{T},\phi^{\star}(\alpha^{\star})\leftarrow\phi^{\star}(\alpha_{T})\}$.